OpenStudy (anonymous):

evaluate the expression -8m-7m^2 m=2

OpenStudy (anonymous):

\[-8m-7m ^{2}\]

OpenStudy (anonymous):

-22m=2 -22/2=-11 your answer is -11

OpenStudy (anonymous):

thanks

OpenStudy (akashdeepdeb):

\[-8m - 7m^2\] Now since m=2, substitute m as 2. \[-8(2) - 7(2)^2\] \[-16-7(4)\] \[-16-28\] \[-44\] Understood? :)

OpenStudy (anonymous):

Yea I'm pretty sure that's right. wait are you saying m=2 or the question was -8m-7m3=2?

OpenStudy (akashdeepdeb):

We have to calculate the value of the expression when m=2. :)